The length of a rectangular driveway is four feet less than five times the width. The area is 672ft^2. Find the width and length of the driveway.
Answer by ikleyn(52781)

.

The area  L*W = 672.


    L = 5W-4


The area  (5W - 4)*W = 672.


    -4W + 5W^2 = 672


    5W^2 - 4W - 672 = 0.


Solve this quadratic equation using the quadratic formula.